ridium-192 is one radioisotope used in brachytherapy, in which a a radioactive source is placed inside a near the tumor while lowering the risk of damage to healthy tissue. Iridium-192 is often used in the head or breast Answer the following three questions (a, b, and c) based on the radioactive decay curve of iridium-192 shown below. Click on the graph and move the mouse over a point to get values for that point Sample remaining 100 90 80 70 60 50 40 30 20 10 10 20 30 40 50 60 70 80 90 100 110 120 130 140 150 160 170 180 190 200 Time (days) a) If the initial sample is 8.00 g, what mass of the original iridium-192 remains after 150 days?
Answer #1

a) After 150days ,25% of the initial mass of the sample will remain, according to the given graph.

Therefore 25% of 8.00g= 2g

b)Original sample =8.00g

Final sample = 2.00g

So how many half-life must have elapsed to get 0.25g

8.00 g...1st half life..... 4.00g...2nd halflife.....2.00g

Therefore after two half lives the final sample becomes 2.00g

hence it is 150/2 = 75 days.

t1/2 = 75days.

c) 1/3 of the sample is nothing but

1/3 * 8= 2.666= 2.67 g

N0 = initial amount = 8.00g

N= final amount = 2.67g

t1/2 = half life= 75 days

t= number of days taken to decay from 8.00g to 2.67g

t= t1/2 * log 1/2 ( N/N0)

t= 75 * log 1/2 ( 2.67 / 8.00)